### v1

# MoPD: Mixture-of-Prompts Distillation for Vision-Language Models

Abstract
Soft prompt learning methods are effective for adapting vision-language models (VLMs) to downstream tasks. Nevertheless, empirical evidence reveals a tendency of existing methods that they overfit seen classes and exhibit degraded performance on unseen classes. This limitation is due to the inherent bias in the training data towards the seen classes. To address this issue, we propose a novel soft prompt learning method, named Mixture-of-Prompts Distillation (MoPD), which can effectively transfer useful knowledge from hard prompts manually hand-crafted (a.k.a. teacher prompts) to the learnable soft prompt (a.k.a. student prompt), thereby enhancing the generalization ability of soft prompts on unseen classes. Moreover, the proposed MoPD method utilizes a gating network that learns to select hard prompts used for prompt distillation. Extensive experiments demonstrate that the proposed MoPD method outperforms state-of-the-art baselines especially on on unseen classes.
Index Terms:
Vision-language models, prompt learning, few-shot learning, prompt distillation.
I Introduction
With large-scale training data of image-text pairs, large pre-trained vision-language models (VLMs) such as CLIP [1], BLIP [2], Flamingo [3], and LLaVA [4] possess a remarkable zero-shot generalization ability and address various downstream vision tasks, such as image classification [1, 5, 6], object detection [7], segmentation [8], visual question answering [2, 3, 9, 10], and visual reasoning [4]. Some soft prompt learning methods, such as CoOp [11] and CoCoOp [12], are built on VLMs. Those methods incorporate learnable prompting vectors into the text encoder or decoder while keeping the pre-trained model fixed, enabling few-shot fine-tuning of VLMs. These learnable prompts act as data-efficient learners and enable large pre-trained VLMs to perform better on a variety of downstream tasks.
Figure 1: Comparison among single-prompt distillation, multi-prompts distillation, and the proposed MoPD method, where single-prompt distillation is to distill knowledge from a single hard prompt to the soft prompt and multi-prompts distillation is to distill knowledge from multiple hard prompts to the soft prompt. ‘H’ denotes the harmonic mean accuracy.
Recently, some works (e.g., ProGrad [13] and KgCoOp [14]) have shown that soft prompt learning methods tend to overfit the base (seen) classes and generalize poorly on the new (unseen) classes, while VLMs with hard prompts (e.g., a photo of a [CLASS], where [CLASS] denotes the placeholder for the class name) can achieve good zero-shot recognition performance on new classes. To utilize the useful knowledge captured in the pre-trained VLMs through hard prompts, ProGrad regulates the gradient direction to mitigate potential conflicts between hard prompts and soft prompts. Similarly, KgCoOp mitigates the risk of knowledge loss by minimizing the embedding discrepancy between soft prompts and hard prompts.
However, the aforementioned methods solely rely on a single hard prompt, without considering the use of multiple hard prompts to enhance the learning of soft prompts. In fact, compared with a single hard prompt, a pool with multiple hard prompts may provide more useful knowledge for prompt-based models and such a pool is commonly available. Besides, knowledge distillation techniques [15, 16, 17, 18] are an effective approach to transfer knowledge, but they are unexplored in existing soft prompt learning methods for VLMs.
Based on the above considerations, in this paper we perform the first investigation on the effectiveness of multi-prompts distillation for VLMs by comparing with single-prompt distillation, where the distillation is from a hard prompt (a.k.a. teacher prompt) to the soft prompt (a.k.a. student prompt). According to the results shown in Fig. 1, we can see that simultaneously distilling the knowledge of multiple hard prompts can achieve better performance when compared with the single hard prompt, which verifies the effectiveness of the use of multiple hard prompts. However, the multi-prompts distillation may be vulnerable to the unrelated and noisy content of hard prompts. Moreover, it is challenging yet beneficial to find an instance-specific teacher prompt for each image due to the varying prompt preferences across different images [12].
To solve these issues, we propose a Mixture-of-Prompts Distillation (MoPD) method. Specifically, the proposed MoPD method utilizes a trainable gating network to select one or more instance-specific hard prompts from a pool of hard prompts, guiding the learning of the soft prompt, where the gating network assigns weights to teacher prompts based on image features during the prompt learning stage. Extensive experiments on 11 datasets demonstrate that the proposed MoPD method outperforms state-of-the-art baseline methods under several settings (i.e., base-to-new, few-shot classification, and domain generalization).
The main contributions of this work are three-fold.
  • We are the first to introduce instance-specific prompt distillation for VLMs to transfer useful knowledge from hard prompts to soft prompts and validate that distilling multiple hard prompts is more effective compared to a single hard prompt.
  • We propose Mixture-of-Prompts Distillation for VLMs, which introduces a gating network to select suitable hard prompts for effectively transferring useful knowledge to the soft prompt.
  • Extensive experimental results demonstrate the superior performance of MoPD. Moreover, our experiments have substantiated the robustness of MoPD to noisy prompts within the hard prompt pool.
II Related Work
II-A Vision-Language Models
In recent years, large pre-trained VLMs [1, 2, 19, 3, 4] have shown great impacts in computer vision. A representative work is CLIP [1], which jointly trains an image encoder built on a Vision Transformer (ViT) [20] or ResNet [21] and a transformer-based text encoder with 400 million image-text pairs via a contrastive objective, showing excellent zero-shot generalization capability. ALIGN [19] is closely related to CLIP, with the difference that ALIGN uses uncleaned training data, so the scale of the collected training data can be larger but noisier. In this paper, we use CLIP as the foundation model.
II-B Prompting
Prompting techniques [22, 23] enable the effective utilization of knowledge in pre-trained models by reformulating the downstream task into a “fill-in-the-blank” cloze test format through prompts. In VLMs, prompts are extensively employed to efficiently adapt the model to various downstream tasks while keeping a small number of training parameters. For instance, CLIP converts a conventional prediction task into the task of associating captions with images, and enables zero-shot transfer to downstream tasks by designing hard prompts like “a photo of a [CLASS]”. However, designing prompts for different downstream tasks is time-consuming and always requires domain expertise, and the designed prompts are not guaranteed to be optimal.
II-C Soft Prompt Learning
To solve the above issue in prompting, soft prompt learning methods are proposed to learn a group of prompting vectors while fixing parameters in the backbone. For example, CoOp [11] utilizes learnable prompting vectors to replace the prompting words like “a photo of a” and adapts to downstream tasks. However, CoOp tends to overfit seen classes, thereby limiting its ability to generalize to unseen classes within the same dataset. To address the issue of weak generalization ability, CoCoOp [12] trains a lightweight neural network to generate input-conditional vectors for each image, collaborating with prompting vectors to dynamically generate prompting vectors. As discussed before, ProGrad and KgCoOp aim to use a hard prompt to guide the learning of the soft prompt in different ways.
To the best of our knowledge, prompt distillation for VLMs remains unexplored. Given the potential of this technique to transfer knowledge, this paper investigates the efficacy of single-prompt distillation and multi-prompt distillation in the context of VLMs.
III Methodology
In this section, we first briefly review CLIP and CoOp, and then introduce the proposed MoPD method in detail.
Figure 2: Overview of the proposed MoPD method.
III-A Preliminaries
CLIP is pre-trained on about 400 million image-text pairs, resulting in its remarkable zero-shot image recognition capability. CLIP comprises two encoders: an image encoder, which is responsible for extracting visual information, and a text encoder, which is used to extract textual information.
Given a set of class names with a total number of $C$ classes, a hard prompt like “a photo of a [$\mathrm{CLASS}$]” is inputted into the text encoder to generate textual embedding $\boldsymbol{t}_{\mathrm{hard}}^{c}$ for the $c$-th class. The visual embedding $\boldsymbol{f}$ is extracted by the image encoder for an image $\boldsymbol{x}$. The prediction probability of image $\boldsymbol{x}$ belonging to class $y$ is formulated as
$p_{\mathrm{hard}}\left(y|\boldsymbol{x}\right)=\frac{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{hard}}^{y},\boldsymbol{f}\right)/\tau\right)}{\sum\nolimits_{c=1}^{C}{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{hard}}^{c},\boldsymbol{f}\right)/\tau\right)}},$ (1)
where $\tau$ is a temperature parameter and $\cos\left(\cdot,\cdot\right)$ denotes the cosine similarity.
However, hard prompts (e.g., “a photo of a [$\mathrm{CLASS}$]”), which are straightforward for humans, may not be optimal for VLMs. To enhance the adaptability of prompts for various downstream tasks, the CoOp method replaces such prompting words with learnable prompting vectors, which are called soft prompts. This soft prompting approach has demonstrated improved performance on multiple downstream datasets compared with using hard prompts [1, 19]. Specifically, the soft prompt consists of learnable prompting vectors of length $M$, where each prompting vector has the same dimension as the word embedding in VLMs, and a class name. Hence, the soft prompt can be defined as $\left\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{M},\left[\mathrm{CLASS}\right]\right\}$. The textual embedding of $c$-th class can be obtained by feeding the soft prompt into the text encoder, and it is denoted by $\boldsymbol{t}_{\mathrm{soft}}^{c}$. Then the prediction probability of image $\boldsymbol{x}$ belonging to class $y$ is formulated as
$p_{\mathrm{soft}}\left(y|\boldsymbol{x}\right)=\frac{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{y},\boldsymbol{f}\right)/\tau\right)}{\sum\nolimits_{c=1}^{C}{\exp\left(\cos\left(\boldsymbol{t}_{\mathrm{soft}}^{c},\boldsymbol{f}\right)/\tau\right)}}.$ (2)
The soft prompt can be learned by minimizing the cross-entropy loss as
$\zeta_{\mathrm{CE}}=-\sum_{(\boldsymbol{x},y)\in\mathcal{D}}{{\ln p_{\mathrm{soft}}\left(y|\boldsymbol{x}\right)}},$ (3)
where $\mathcal{D}$ denotes the training set. Note that the image encoder and text encoder are frozen during the process of soft prompt learning.
III-B Mixture-of-Prompts Distillation
Besides a frozen image encoder and a frozen text encoder for extracting visual and textual information as well as a pool of hard prompts, MoPD consists of a soft prompt and a trainable gating network to select instance-specific hard prompts for each image to guide the process of soft prompt learning. The overall framework of MoPD is shown in Fig. 2.
III-B1 Single-Prompt Distillation (SiPD)
Here we first introduce how to perform distillation to guide the learning of a soft prompt based on a hard prompt, which will set the stage for the following introduction of MoPD.
Knowledge distillation aims at transferring the knowledge from a teacher model to a student model. Similarly, we introduce prompt distillation as a technique to enhance the generalization ability of soft prompts based on hard prompts. In prompt distillation, a hard prompt serves as the teacher prompt, guiding the learning of the soft prompt, which acts as the student prompt. Considering Eqs. (1) and (2), the prediction probability distributions of the hard and soft prompts can be formulated as $p_{\mathrm{hard}}(\cdot|\boldsymbol{x})$ and $p_{\mathrm{soft}}(\cdot|\boldsymbol{x})$, respectively. Then the prompt distillation loss is formulated as
$\zeta_{\mathrm{PD}}=\sum_{\boldsymbol{x}\in\mathcal{D}}{{\mathrm{KL}\left(p_{\mathrm{soft}}\left(\cdot|\boldsymbol{x}\right),p_{\mathrm{hard}}\left(\cdot|\boldsymbol{x}\right)\right)}},$ (4)
where $\mathrm{KL}\left(\cdot,\cdot\right)$ denotes the KL divergence. By minimizing $\zeta_{\mathrm{PD}}$, we can transfer useful knowledge contained in the teacher prompt to the student prompt.
III-B2 Mixture-of-Prompts Distillation (MoPD)
Usually there may exist significant variations among images within the same dataset [24] and hence a single hard prompt is insufficient to accommodate the preferences of all images. To alleviate this issue, it is intuitive to utilize multiple hard prompts to fit such preferences and a straightforward way is to randomly select hard prompts from a hard prompt pool. However, this method still does not take the individual preferences of each image into consideration and cannot filter out potentially noisy hard prompts from the pool.
To find a better way to distill the knowledge from multiple hard prompts, inspired by mixture-of-experts [25], the proposed MoPD method additionally introduces a gating network $G$ to select suitable and instance-specific hard prompts for each image to guide the learning of the soft prompt. For simplicity, the gating network $G$ has only one fully connected layer with a trainable weight matrix $\boldsymbol{W}_{g}$ and adopts the softmax function as the activation function. The input of the gating network $G$ is the visual embedding $\boldsymbol{f}$ produced by the image encoder. Then the gating network can be formulated as
$G(\boldsymbol{f})=\mathrm{Softmax}(\mathrm{KeepTop}(\boldsymbol{f}\cdot\boldsymbol{W}_{g},T)),$ (5)
where $\mathrm{Softmax}(\cdot)$ denotes the softmax function and $\mathrm{KeepTop}(\cdot,\cdot)$ is defined as
$\mathrm{KeepTop}(u,T)=\begin{cases}u_{i},&\mathrm{if}\;u_{i}\;\mathrm{is}\;\mathrm{in}\;\mathrm{top}\;T\;\mathrm{elements}\;\mathrm{of}\;u\\ -\infty,&\mathrm{otherwise}\end{cases}.$ (6)
The output of the gating network represents the preference of the image towards all the hard prompts. Since not all the hard prompts are informative, we only choose the top-$T$ hard prompts via the $\mathrm{KeepTop}(\cdot,\cdot)$ operator.
Once the hard prompts are selected, we formulate a mixture-of-prompts distillation loss to distill knowledge from these hard prompts (i.e., teacher prompts) to the soft prompt (i.e., student prompt) as
$\zeta_{\mathrm{MPD}}=\sum_{t=1}^{H}{\sum_{\boldsymbol{x}\in\mathcal{D}}{{G\left(\boldsymbol{f}\right)_{t}\mathrm{KL}\left(p_{\mathrm{soft}}\left(\cdot|\boldsymbol{x}\right),p_{\mathrm{tea},t}\left(\cdot|\boldsymbol{x}\right)\right)}}},$ (7)
where $p_{\mathrm{tea},t}\left(\cdot|\boldsymbol{x}\right)$ denotes the prediction probability distribution of the $t$-th teacher prompt for all the classes based on Eq. (1), $G(\boldsymbol{f})_{t}$ denotes the weight corresponding to $t$-th teacher prompt, and $H$ denotes the total number of hard prompts in the pool.
To enhance the gating network’s ability to autonomously select the most suitable teacher prompts for downstream tasks and effectively mitigate interference from noisy prompts, we incorporate the prediction probability of teacher prompts into the loss function to guide the learning process of this gating network. Specifically, by encouraging the gating network to assign higher weights to teacher prompts that have higher prediction probabilities for the correct label, the gating network learns to prioritize the most informative and relevant teacher prompts for the downstream task, while suppressing the influence of noisy or less relevant prompts. Therefore, we propose the mixture-of-prompts selection loss as
$\zeta_{\mathrm{MPS}}=-\sum_{t=1}^{H}{\sum_{(\boldsymbol{x},y)\in\mathcal{D}}{{G\left(\boldsymbol{f}\right)_{t}\ln p_{\mathrm{tea},t}\left(y|\boldsymbol{x}\right)}}}.$ (8)
By combining the above considerations together, the overall objective of the MoPD method is formulated as
$\zeta=\alpha\zeta_{\mathrm{CE}}+\left(1-\alpha\right)\zeta_{\mathrm{MPD}}+\beta\zeta_{\mathrm{MPS}},$ (9)
where $\alpha$ and $\beta$ are trade-off parameters. Here $\zeta_{\mathrm{CE}}$ and $\zeta_{\mathrm{MPD}}$ are used to learn the soft prompt, while $\zeta_{\mathrm{MPD}}$ and $\zeta_{\mathrm{MPS}}$ help learn the gating network.
Note that the gating network is only used in the training stage. After it is learned in the training process, it will not be used in the inference stage, making the inference process of MoPD identical to other soft prompt learning methods.
IV Experiments
In this section, we evaluate MoPD on three settings: (1) base-to-new generalization, (2) few-shot classification, (3) domain generalization. We also analyze the effect of each module and parameter through ablation study and parameter analysis. We further analyze the impact of noisy prompts.
Datasets
Following [11, 12, 13, 14], the base-to-new generalization and few-shot image classification are conducted on 11 image classification datasets, including ImageNet [26] and Caltech101 [27] for generic object classification, OxfordPets [28], StanfordCars [29], Flowers102 [30], Food101 [31], and FGVCAircraft [32] for fine-grained visual categorization, SUN397 [33] for scene recognition, DTD [34] for texture classification, EuroSAT [35] for satellite image classification, and UCF101 [36] for action recognition. Besides, the domain generalization is conducted on ImageNet and its variants, including ImageNetV2 [37], ImageNet-Sketch [38], ImageNet-A [39], and ImageNet-R [40].
Training Details
Our implementation is based on the official implementation of CoOp [11] and KgCoOp [14]. All experiments are conducted with ViT-B/16 as the image encoder and Transformer [41] as the text encoder. The gating network is a single-layer fully-connected neural network with the softmax function as the activation function. The length of prompting vectors is fixed to 4, and they are initialized with the template “a photo of a”. The training schedule and data augmentation settings are identical to those of CoOp and KgCoOp. All experimental results are averaged over three seeds. In Eq. (9), $\alpha$ is set to 0.8 by default, except that $\alpha$ is set to 0.5 on Food101, EuroSAT, and UCF101. $\beta$ is set to 0.0005 by default, except that $\beta$ is set to 1.0 on StanfordCars and DTD. The number of selected teacher prompts (i.e., $T$) in the gating network is set to 3 on ImageNet, Caltech101, OxfordPets, StanfordCars, and Flowers102, while for the rest datasets, $T$ is set to 2. The hard prompt pool consisting of task-related hard prompts is simply designed through synonym substitution and sentence transformation on templates of CLIP and CoOp (see Supplementary Material for a full list), and the total number of hard prompts (i.e., $H$) in the pool is fixed to 12. Similar to [12], we utilize the harmonic mean accuracy $\mathrm{H}=2/\left(1/\mathrm{acc}_{\mathrm{Base}}+1/\mathrm{acc}_{\mathrm{New}}\right)$ as the metric to quantify the trade-off in generalization, where $\mathrm{acc}_{\mathrm{Base}}$ and $\mathrm{acc}_{\mathrm{New}}$ denote the accuracy of the model on the base and new classes, respectively. A high value of H indicates that the corresponding model performs well on both base and new classes simultaneously. We use PyTorch to implement all experiments on NVIDIA GeForce RTX 3090.
Baselines
We compare with several state-of-the-art soft prompt learning methods, including CoOp [11], CoCoOp [12], ProGrad [13], and KgCoOp [14]. CLIP [1] is also in comparison to compare with the generalization performance of hard prompts.
IV-A Results on Base-to-New Generalization
We follow the base-to-new generalization setting of CoCoOp, where all classes are split into disjointed base and new groups and all the methods in comparison are trained on the base classes and evaluated on base and new classes. The experimental results shown in Table I reveal the following observations.
TABLE I: Performance of various methods under the base-to-new generalization setting. All methods are trained with 16 instances per base class. ‘H’ denotes the harmonic mean accuracy.
Dataset Set CLIP CoOp CoCoOp ProGrad KgCoOp MoPD
Average Base 69.34 82.64 80.47 82.48 80.73 81.40
New 74.22 68.00 71.69 70.72 73.61 74.69
H 71.70 74.61 75.83 76.15 77.00 77.90
ImageNet Base 72.43 76.46 75.98 77.02 75.83 76.87
New 68.14 66.31 70.43 66.66 69.96 70.43
H 70.22 71.02 73.10 71.46 72.78 73.51
Caltech 101 Base 96.84 98.11 97.96 98.02 97.72 98.07
New 94.00 93.52 93.81 93.89 94.39 94.90
H 95.40 95.76 95.84 95.91 96.03 96.46
Oxford Pets Base 91.17 94.24 95.20 95.07 94.65 95.17
New 97.26 96.66 97.67 97.63 97.76 97.47
H 94.12 95.43 96.43 96.33 96.18 96.31
Stanford Cars Base 63.37 76.20 70.49 77.68 71.76 75.63
New 74.89 69.14 73.59 68.63 75.04 73.10
H 68.65 72.49 72.01 72.88 73.36 74.34
Flowers 102 Base 72.08 97.63 94.87 95.54 95.00 95.80
New 77.80 69.55 71.75 71.87 74.73 74.43
H 74.83 81.23 81.71 82.03 83.65 83.77
Food101 Base 90.10 89.44 90.70 90.37 90.50 90.77
New 91.22 87.50 91.29 89.59 91.70 91.67
H 90.66 88.46 90.99 89.98 91.09 91.22
FGVC Aircraft Base 27.19 39.24 33.41 40.54 36.21 37.43
New 36.29 30.49 23.71 27.57 33.55 35.13
H 31.09 34.30 27.74 32.82 34.83 36.24
SUN397 Base 69.36 80.85 79.74 81.26 80.29 81.13
New 75.35 68.34 76.86 74.17 76.53 76.47
H 72.23 74.07 78.27 77.55 78.36 78.73
DTD Base 53.24 80.17 77.01 77.35 77.55 77.27
New 59.90 47.54 56.00 52.35 54.99 57.47
H 56.37 59.68 64.85 62.45 64.35 65.92
EuroSAT Base 56.48 91.54 87.49 90.11 85.64 85.77
New 64.05 54.44 60.04 60.89 64.34 72.50
H 60.03 68.27 71.21 72.67 73.48 78.58
UCF101 Base 70.53 85.14 82.33 84.33 82.89 81.53
New 77.50 64.47 73.45 74.94 76.67 78.07
H 73.85 73.37 77.64 79.35 79.65 79.76
  1. 1.
    In terms of the harmonic mean accuracy, MoPD outperforms all baseline methods, showcasing its excellent generalization capability on the base-to-new setting. Specifically, MoPD attains the highest harmonic mean accuracy in 10 out of 11 datasets.
  2. 2.
    In terms of the average performance on the new classes, MoPD surpasses all baseline methods. MoPD exhibits a significant performance improvement over soft prompt learning methods and outperforms CLIP on 7 out of 11 datasets. This demonstrates that the selected teacher prompts can effectively guide the soft prompt in acquiring useful knowledge and enhancing its generalization capability.
  3. 3.
    In terms of the average performance on the base classes, CoOp and ProGrad perform better among all methods since they may overfit the base classes, leading to poor performance on new classes. MoPD performs slightly inferior to CoOp and ProGrad on the base classes, but substantially outperforms them on the new classes, while MoPD outperforms CoCoOp and KgCoOp on both the base and new classes.
IV-B Results on Few-shot Classification
In the few-shot classification setting, each method is trained on all the classes of a dataset with few-shot labeled samples and evaluated on the rest data of this dataset. Following [11], we perform few-shot classification with different numbers of shots, specifically 1, 2, 4, 8, and 16 shots. The performance of all the methods on the 11 datasets is shown in Fig. 3. According to the results, MoPD achieves the best average performance among all baseline methods. Those results further demonstrate the good generalization ability of the proposed MoPD method when the samples of downstream tasks are limited.
Figure 3: Performance of various methods under the few-shot classification setting on the 11 datasets.
IV-C Results on Domain Generalization
Generalisation to out-of-distribution data is a necessary capability for VLMs. The domain generalization setting is employed to evaluate the generalization ability of a model trained on a source domain to a target domain that shares the same classes as the source domain but exhibits a different data distribution. As with all baseline methods (except zero-shot CLIP), we perform soft prompt learning on the source domain (i.e., ImageNet) and evaluate the model on four different target domains (i.e., ImageNetV2, ImageNet-Sketch, Imagenet-A, and ImageNet-R).
According to results shown in Table II, we can observe that MoPD outperforms all baseline methods except ProGrad in the source domain, and MoPD performs better than all baseline methods on ImageNet-Sketch and ImageNet-A. In terms of average performance on all the four target domains, MoPD performs the best among all baseline methods, which confirms that the soft prompts learned in MoPD are more domain-generalizable.
TABLE II: Performance of various methods under the domain generalization setting. All methods use 16-shot samples from each of the 1000 classes on ImageNet except CLIP. The best and second-best results are marked in bold and underlined, respectively.
Source Target
ImageNet -V2 -Sketch -A -R Avg.
CLIP 66.73 60.83 46.15 47.77 73.96 57.17
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.90
ProGrad 72.24 64.73 47.61 49.39 74.58 59.07
KgCoOp 71.20 64.10 48.97 50.69 76.70 60.11
MoPD 71.77 64.70 49.10 50.77 76.60 60.29
IV-D Ablation Study
We conduct the ablation study to demonstrate the effect of mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and the selection of gating network under the base-to-new generalization setting.
The results are listed in Table III, where the baseline is CoOp, SiPD built on CoOp is formulated in Section III-B, MoPD-R denotes a variant of MoPD by randomly selecting teacher prompts instead of using the gating network, and MoPD (w/o $\zeta_{\mathrm{MPS}}$) denotes MoPD without minimizing $\zeta_{\mathrm{MPS}}$.
[…]
TABLE III: Ablation study for the mixture-of-prompts distillation loss $\zeta_{\mathrm{MPD}}$, mixture-of-prompts selection loss $\zeta_{\mathrm{MPS}}$, and gating network.
Dataset Set Baseline SiPD MoPD-R MoPD MoPD
(CoOp) (w/o $\zeta_{\mathrm{MPS}}$)
Average Base 82.64 81.11 81.39 81.54 81.40
New 68.00 73.35 73.52 74.19 74.69
H 74.61 77.04 77.26 77.69 77.90
Figure 4: The average performance of all datasets achieved by MoPD for hyperparameters (i.e., $\alpha$, $\beta$, $T$, and $H$) under the base-to-new generalization setting.
IV-E Parameter Analysis
In this section, we analyze the effect of several hyperparameters (i.e., $\alpha$, $\beta$, $T$, and $H$) in MoPD. The average performance achieved by MoPD on all datasets under the base-to-new generalization setting is shown in Fig. 4.
For the trade-off parameter $\alpha$, according to Fig. 4a, we can see that as $\alpha$ increases, the accuracy on base classes increases, while that on new classes increases and then decreases. In other words, increasing $\alpha$ will increase the weight of $\zeta_{\mathrm{CE}}$, allowing the soft prompt to fit better on base classes, but will deprive its generality on new classes. On the other hand, decreasing $\alpha$ to increase the weight of $\zeta_{\mathrm{MPD}}$ will enable the soft prompt to generalize better on new classes. This suggests that it is necessary to choose a medium value for $\alpha$, which is usually set to 0.5 or 0.8 to achieve a better harmonic mean accuracy in our experiments.
For the trade-off parameter $\beta$, Fig. 4b illustrates that $\beta$ within some range (e.g., 0.00005 to 10) has little impact on base classes. In terms of the performance on new classes and their harmonic mean accuracy, the performance shows a trend of initially increasing and then decreasing within some range of $\beta$ (e.g., 0.00005 to 0.5). Those results suggest that $\beta$ could take values around 0.0005.
For the number of selected teacher prompts $T$, it is evident from Fig. 4c that as $T$ increases, accuracy on base classes increases slightly, and the accuracy on new classes increases first and then decreases. A large $T$ will weaken the effect of each teacher prompt, leading to inferior performance. When $T=2$, the average H over all datasets is the highest. It is important to note that when $T=1$, the teacher prompt for the same image may differ across different training iterations. Consequently, the image also benefits from guidance provided by multiple teacher prompts. This is a key factor contributing to MoPD’s strong performance, where MoPD surpasses SiPD by a considerable margin even when $T=1$.
For the parameter $H$, we expand the range of the number of hard prompts in pool from 1 to 36 and as shown in Fig. 4d, the performance of MoPD generally improves as the number of prompts increases from 1 to 12. However, further augmentation of the prompt quantity yields no significant performance enhancement. This observation suggests that continually increasing the number of prompts does not confer substantial benefits.
IV-F Further Analysis
To investigate the robustness of MoPD to noisy prompts, we add noisy prompts into the hard prompt pool and compare MoPD with MoPD-R under the base-to-new generalization setting on the UCF101 and DTD datasets, where noisy prompts are randomly generated based on grammatically plausible sentence templates with incoherent words generated by ChatGPT [42, 43], such as “In the whispering forest, a [CLASS] dreams of stars” (more examples can be found in Supplementary Material).
According to the results shown in Table IV, when the hard prompt pool is full of noisy prompts (i.e., 24N and 12N), the performance of both MoPD and MoPD-R is clearly worse than MoPD and MoPD-R with 12T. Nevertheless, even in situations where the hard prompt pool is exclusively composed of noisy prompts, MoPD is capable of discerning relatively superior teacher prompts, resulting in higher harmonic mean accuracy compared to MoPD-R. When the hard prompt pool is a mixture of noisy prompts and task-related prompts, the higher the proportion of noisy prompts in the prompt pool, the worse the performance (in terms of base, new, and harmonic mean accuracy) of MoPD-R. However, the performance of MoPD with 12T shows almost no difference when compared to MoPD with 12T+12N and MoPD with 12T+24N. This is because the gating network can help select suitable teacher prompts. Those experimental results demonstrate the robustness of MoPD against noisy prompts.
Furthermore, we investigate the impact of load balancing [25] on MoPD in the Supplementary Material. The results indicate that the performance comparison between the balanced and unbalanced versions is negligible. Consequently, in our experiments, we present results for the basic MoPD implementation without load balancing. Additional details and comprehensive analyses of this aspect can be found in the Supplementary Material.
TABLE IV: Robustness of MoPD to noisy prompts under the base-to-new generalization setting on the UCF101 and DTD datasets. ‘N’ denotes noisy prompts, ‘T’ denotes task-related prompts, and ‘+’ denotes the mixture of the two types of prompts.
UCF101 MoPD 24N 12N 12T+24N 12T+12N 12T
Base 79.03 79.47 81.57 81.60 81.53
New 74.77 74.30 77.93 78.03 78.07
H 76.84 76.80 79.71 79.78 79.76
MoPD-R 24N 12N 12T+24N 12T+12N 12T
Base 77.93 77.77 79.40 80.33 81.60
New 73.97 74.10 75.30 77.50 77.70
H 75.90 75.89 77.30 78.89 79.60
DTD MoPD 24N 12N 12T+24N 12T+12N 12T
Base 78.20 78.33 77.53 77.63 77.27
New 55.63 55.60 57.50 57.63 57.47
H 65.01 65.04 66.03 66.15 65.92
MoPD-R 24N 12N 12T+24N 12T+12N 12T
Base 78.67 78.07 78.73 78.50 78.27
New 53.80 54.60 54.03 54.80 55.57
H 63.90 64.26 64.08 64.54 64.99
V Conclusion
In this paper, we propose a soft prompt learning method for VLMs named Mixture-of-Prompts Distillation (MoPD) to enhance the generability of soft prompts on unseen classes. MoPD leverages a gating network to select suitable instance-specific teacher prompts for prompt distillation, guiding soft prompts to acquire useful knowledge. Extensive experiments conducted on 11 benchmark datasets demonstrate the effectiveness of MoPD. Notably, MoPD exhibits superior generalization performance on unseen classes, surpassing even CLIP with hard prompts.
Acknowledgments
This work is supported by National Key R&D Program of China (No. 2022ZD0160300) and NSFC key grant 62136005.
Appendix A Implementation Details
A-A Hard Prompt Pool
The hard prompt pool consisting of task-related hard prompts is simply designed through synonym substitution and sentence transformation on prompt templates of CLIP [1] and CoOp [11]. Unless otherwise specified, both MoPD and MoPD-R methods utilize the following hard prompt pool in the experiments. SiPD performs prompt distillation by employing the first prompt from the hard prompt pool.
For ImageNet, Caltech101, StanfordCars, and SUN397, the hard prompt pool is
‘‘a photo of a [CLASS].’’
‘‘a picture of a [CLASS].’’
‘‘an image of a [CLASS].’’
‘‘a photo, classified as [CLASS].’’
‘‘a photo, categorized as [CLASS].’’
‘‘a photo, labeled as [CLASS].’’
‘‘a picture, classified as [CLASS].’’
‘‘a picture, categorized as [CLASS].’’
‘‘a picture, labeled as [CLASS].’’
‘‘an image, classified as [CLASS].’’
‘‘an image, categorized as [CLASS].’’
‘‘an image, labeled as [CLASS].’’
For Food101, the hard prompt pool is
‘‘a photo of [CLASS], a type of food.’’
‘‘a picture of [CLASS], a type of food.’’
‘‘an image of [CLASS], a type of food.’’
‘‘a photo of food, classified as [CLASS].’’
‘‘a photo of food, categorized as
  [CLASS].’’
‘‘a photo of food, labeled as [CLASS].’’
‘‘a picture of food, classified as
  [CLASS].’’
‘‘a picture of food, categorized as
  [CLASS].’’
‘‘a picture of food, labeled as [CLASS].’’
‘‘an image of food, classified as
  [CLASS].’’
‘‘an image of food, categorized as
  [CLASS].’’
‘‘an image of food, labeled as [CLASS].’’
For OxfordPets, Flowers102, and FGVCAircraft, the hard prompt pool is
‘‘a photo of a [CLASS], a type of {}.’’
‘‘a picture of a [CLASS], a type of {}.’’
‘‘an image of a [CLASS], a type of {}.’’
‘‘a photo of a {}, classified as [CLASS].’’
‘‘a photo of a {}, categorized as
  [CLASS].’’
‘‘a photo of a {}, labeled as [CLASS].’’
‘‘a picture of a {}, classified as
  [CLASS].’’
‘‘a picture of a {}, categorized as
  [CLASS].’’
‘‘a picture of a {}, labeled as [CLASS].’’
‘‘an image of a {}, classified as
  [CLASS].’’
‘‘an image of a {}, categorized as
  [CLASS].’’
‘‘an image of a {}, labeled as [CLASS].’’
For DTD, the hard prompt pool is
‘‘a photo of [CLASS], a type of texture.’’
‘‘a picture of [CLASS], a type of
  texture.’’
‘‘an image of [CLASS], a type of texture.’’
‘‘a photo of a texture, classified
  as [CLASS].’’
‘‘a photo of a texture, categorized
  as [CLASS].’’
‘‘a photo of a texture, labeled
  as [CLASS].’’
‘‘a picture of a texture, classified
  as [CLASS].’’
‘‘a picture of a texture, categorized
  as [CLASS].’’
‘‘a picture of a texture, labeled
  as [CLASS].’’
‘‘an image of a texture, classified
  as [CLASS].’’
‘‘an image of a texture, categorized
  as [CLASS].’’
‘‘an image of a texture, labeled
  as [CLASS].’’
For EuroSAT, the hard prompt pool is
‘‘a centered satellite photo of
  [CLASS].’’
‘‘a centered satellite picture of
  [CLASS].’’
‘‘a centered satellite image of
  [CLASS].’’
‘‘a centered satellite photo,
  classified as [CLASS].’’
‘‘a centered satellite photo,
  categorized as [CLASS].’’
‘‘a centered satellite photo,
  labeled as [CLASS].’’
‘‘a centered satellite picture,
  classified as [CLASS].’’
‘‘a centered satellite picture,
  categorized as [CLASS].’’
‘‘a centered satellite picture,
  labeled as [CLASS].’’
‘‘a centered satellite image,
  classified as [CLASS].’’
‘‘a centered satellite image,
  categorized as [CLASS].’’
‘‘a centered satellite image,
  labeled as [CLASS].’’
For UCF101, the hard prompt pool is
‘‘a photo of a person doing [CLASS].’’
‘‘a picture of a person doing [CLASS].’’
‘‘an image of a person doing [CLASS].’’
‘‘a photo of a person doing, classified
  as [CLASS].’’
‘‘a photo of a person doing, categorized
  as [CLASS].’’
‘‘a photo of a person doing, labeled
  as [CLASS].’’
‘‘a picture of a person doing, classified
  as [CLASS].’’
‘‘a picture of a person doing, categorized
  as [CLASS].’’
‘‘a picture of a person doing, labeled
  as [CLASS].’’
‘‘an image of a person doing, classified
  as [CLASS].’’
‘‘an image of a person doing, categorized
  as [CLASS].’’
‘‘an image of a person doing, labeled
  as [CLASS].’’
[…]
A-B Noisy Hard Prompt Pool
The term ”noisy prompts” refers to prompts that are unrelated to the downstream task. Taking inspiration from the generation of random prompts by [44], we employ a similar approach to generate noisy prompts. These noisy prompts are random, grammatically plausible sentence templates with incoherent words generated by ChatGPT [42, 43]. We prompt the language model with the input: “Please generate 12 random prompt templates by sampling grammatically plausible random sentences that contain incoherent words, with length between 5 and 20 words, such as “a photo of a [CLASS]”, where [CLASS] is the class name.” The noisy prompt pool of size 24 is shown as follow.
‘‘A collection of [CLASS] floating in the
  sky.’’
‘‘[CLASS] wrapped in a mysterious bubble.’’
‘‘An unexpected encounter with a [CLASS].’’
‘‘A [CLASS] hidden among the stars.’’
‘‘A vibrant dance of the [CLASS].’’
‘‘The laughter of a [CLASS] echoing through
  the forest.’’
‘‘A [CLASS] emerging from the depths of the
  ocean.’’
‘‘A journey through the pages of a
  [CLASS].’’
‘‘The whispers of a [CLASS] in the wind.’’
‘‘A [CLASS] perched on the edge of a
  cliff.’’
‘‘A [CLASS] illuminated by the moonlight.’’
‘‘The song of a [CLASS] resonating in the
  mountains.’’
‘‘Under the luminous sky, a [CLASS] dances
  with shadows.’’
‘‘In the whispering forest, a [CLASS]
  dreams of stars.’’
‘‘A [CLASS] floats in a teacup filled
  with stormy clouds.’’
‘‘With a melody of forgotten songs, a
  [CLASS] paints the sunrise.’’
‘‘A [CLASS] juggles moonbeams at the edge
  of twilight.’’
‘‘Inside a clockwork, a [CLASS] unravels
  the fabric of time.’’
‘‘Beneath the rainbow waterfall, a [CLASS]
  plays a silent tune.’’
‘‘A [CLASS] weaves dreams into the fabric
  of the cosmos.’’
‘‘In the heart of an ancient book, a
  [CLASS] discovers hidden worlds.’’
‘‘A [CLASS] dances on the petals of an
  upside-down flower.’’
‘‘Amongst the whispers of the wind, a
  [CLASS] listens to the earth’s
  heartbeat.’’
‘‘A [CLASS] rides the waves of a sunbeam
  across the ocean of eternity.’’
Note that in Table IV, the notation “12N” denotes that MoPD and MoPD-R are trained utilizing the first 12 noisy prompts listed in the table. While the notation ”24N” denotes training with all 24 noisy prompts in the table.
Appendix B Experimental Details
B-A Details for Datasets
The detailed statistics of the 11 datasets for the base-to-new generalization setting are shown in Table A5, where the soft prompts are learned with 16 instances per base class.
TABLE A5: The detailed statistics of datasets under the base-to-new generalization setting.
Datasets Number of classes Number of samples
Base New Train Test-Base Test-New
ImageNet 500 500 8000 16402 16300
Caltech101 50 50 800 1549 916
OxfordPets 19 18 304 1881 1788
StanfordCars 98 98 1568 4002 4039
Flowers102 51 51 816 1053 1410
Food101 51 50 816 15300 15000
FGVCAircraft 50 50 800 1666 1667
SUN397 199 198 3184 9950 9900
DTD 24 23 384 864 828
EuroSAT 5 5 80 4200 3900
UCF101 51 50 816 1934 1849
Appendix C Additional Results on Load Balance
Inspired by SMoP [45], we adopt load balance for exploring diverse teacher prompts by adding a scaled Gaussian noise $\delta\sim\mathcal{N}\left(0,1\right)$, which decreases within the training process, onto $(\boldsymbol{f}\cdot\boldsymbol{W}_{g})$ in Eq. (5) in Sec. III-B. Consequently, Eq. (5) is modified as
$G(\boldsymbol{f})=\mathrm{Softmax}(\mathrm{KeepTop}((\boldsymbol{f}\cdot\boldsymbol{W}_{g}\cdot(\mathbf{I}+\delta\cdot p))),T))$ (10)
where $\mathbf{I}$ is the identity matrix, and $p\in\left[0,1\right]$ is the ratio of current to maximum iterations.
The results for the base-to-new generalization setting is shown in Table A6, where MoPD-L represents a variant of MoPD that uses load balance, the task-related hard prompt pool means that its hard prompts are simply designed by synonym substitution and sentence transformation on CLIP’s prompt templates, and the mixed hard prompt pool means the mixture of task-related and task-unrelated hard prompt pool.
TABLE A6: The effect of load balance under the base-to-new generalization setting on all datasets. The numbers in red and blue denote the best performance from different prompt pools, respectively.
Dataset Set Task-related hard prompt pool Mixed hard prompt pool
MoPD MoPD MoPD MoPD MoPD MoPD
-R -L -R -L
Average Base 81.39 81.28 81.40 80.74 81.09 81.04
New 73.52 74.64 74.69 73.92 74.76 74.52
H 77.26 77.82 77.90 77.18 77.80 77.64
ImageNet Base 76.87 76.83 76.87 76.97 77.00 76.93
New 70.10 70.10 70.43 70.17 69.87 70.30
H 73.33 73.31 73.51 73.41 73.26 73.47
Caltech 101 Base 98.07 97.90 98.07 97.90 97.83 97.97
New 94.50 95.50 94.90 95.27 94.73 94.57
H 96.25 96.69 96.46 96.57 96.26 96.24
Oxford Pets Base 95.13 94.80 95.17 95.13 95.30 95.33
New 97.40 97.17 97.47 97.27 96.83 97.57
H 96.25 95.97 96.31 96.19 96.06 96.44
Stanford Cars Base 75.70 75.27 75.63 75.33 75.50 75.83
New 72.70 73.57 73.10 70.33 72.53 72.40
H 74.17 74.41 74.34 72.74 73.99 74.08
Flowers 102 Base 96.30 96.13 95.80 95.63 96.00 95.43
New 72.90 73.27 74.43 72.03 73.53 74.23
H 82.98 83.16 83.77 82.17 83.28 83.51
Food101 Base 90.60 90.80 90.77 90.57 90.73 90.70
New 91.60 91.67 91.67 91.73 91.97 91.97
H 91.10 91.23 91.22 91.15 91.35 91.33
FGVC Aircraft Base 36.50 36.47 37.43 36.93 35.03 37.90
New 34.70 36.50 35.13 34.73 35.83 35.07
H 35.58 36.48 36.24 35.80 35.43 36.43
SUN397 Base 80.87 81.27 81.13 80.87 80.77 80.90
New 76.17 76.57 76.47 76.00 76.83 76.43
H 78.45 78.85 78.73 78.36 78.75 78.60
DTD Base 78.27 78.43 77.27 77.57 79.00 75.13
New 55.57 56.60 57.47 57.33 57.80 59.33
H 64.99 65.75 65.92 65.93 66.76 66.30
EuroSAT Base 85.40 84.83 85.77 80.07 83.80 83.77
New 65.37 71.73 72.50 71.00 74.80 69.33
H 74.05 77.73 78.58 75.26 79.04 75.87
UCF101 Base 81.60 81.37 81.53 81.13 81.00 81.53
New 77.70 78.37 78.07 77.27 77.67 78.50
H 79.60 79.84 79.76 79.15 79.30 79.99
With the task-related hard prompt pool, compared to MoPD-R and MoPD-L, MoPD achieves the best average performance (in terms of base, new, and harmonic mean accuracy). MoPD-L is slightly inferior to MoPD but also achieves good performance. With the mixed hard prompt pool, MoPD-L achieves slightly better performance than MoPD. Due to the negligible difference in average performance between MoPD and MoPD-L, we use the MoPD method in all the experiments.